# Edge Detectors Can Make Deep Convolutional Neural Networks More Robust

### v1

@@ 1 Introduction
It is well known that deep convolutional neural networks (DCNN for short) can be fooled by examples with small perturbations Goodfellow2015ICLR ; Szegedy2014ICLR ; Tao2021IJCAI ; Zhang2019TNNLS ; Serban2020CSUR , which brings the potential hazards when applying DCNNs to the safety-critical applications, e.g., autonomous driving, airport security, and industry automation. Therefore, it is of great significance to improve the robustness of DCNNs. Fig. 1 shows several clean examples and their adversarial counterparts. Adversarial examples (AE for short) in Fig. 1(b) are with small noises which can induce DCNNs to make wrong decisions. However, these noises can be filtered by human eyes easily. Researches show that the principal way human beings recognize the objects is largely relying on the shape features Davitt2014JEP ; Sigurdardottir2014JEP , and that’s why human beings are not susceptible to the noise imposed in Fig. 1(b). Inspired by this, it is natural to ask, is it possible to make the DCNNs more robust by learning the shape-like features?
@@
Fig. 2 shows the thresholded edge images of Fig. 1. From the figure, it can be observed that the thresholded edge images for Fig. 1(a) and Fig. 1(b) are almost the same. It prompts us that, enhancing the robustness of DCNNs can be achieved by using the binary edge features which can be seen as a kind of shape features. To this end, in this paper, four learnable edge detectors are designed and taken as layer kernels, which can be used to extract horizontal, vertical, positive diagonal, and negative diagonal edge features, respectively Kittler1983IVC ; Xiao2022JOS ; Yasir2022Science . And based on the edge detectors, a binary edge feature branch (BEFB for short) is proposed, which is staked by multiple Sobel layers and one threshold layer. The Sobel layers employ the edge detectors as kernels, and the threshold layer turns the output of the last Sobel layer to the binary features. The binary features concatenated with the texture features learnt by any popular backbone are then fed into fully connected layers for classification. To deal with the zero gradient of threshold layer, which makes the weights in BEFB unable to update using the chain rule, STE technique Bengio2013Arxiv ; Yang2022ICML ; Vanderschueren2023WACV is employed. In addition, to take obfuscated gradient effect raised by Athalye2018ICML ; Yue2023USENIX ; Huang2022MSML into consideration, zero gradient or one center-translated sigmoid activation function are also used for generating AEs. In the experiments, we integrate BEFB into VGG16 Simonyan2014ICLR and ResNet34 He2016CVPR ; He2020TNNLS , respectively, and conduct experiments on multiple datasets. The results demonstrate BEFB has no side effects on model training, and BEFB integrated models can achieve better accuracy under FGSM Goodfellow2015ICLR , PGD Kurakin2017ICLR , and C&W Carlini2017SP attacks compared to the original ones. Furthermore, we combine the BEFB integrated models with two popular robustness enhancing techniques–AT(abbreviation for adversarial training) Madry2018ICLR and PCL(abbreviation for prototype conformity loss) Mustafa2021TPAMI , respectively, and find BEFB integrated models can achieve better accuracy than the original models as well.
@@ 2 Related Work
Robustness Enhancing with AT. AT improves the robustness of the DCNNs by generating AEs as training samples during optimization. Tsipras et al. Tsipras2019ICLR found there exists a tradeoff between robustness and standard accuracy of the models generated by AT, due to the robust classifiers learning a different feature representation compared to the clean classifiers. Madry et al. Madry2018ICLR proposed an approximated solution framework to the optimization problems of AT, and found PGD-based AT can produce models defending themselves against the first-order adversaries. Kannan et al. Kannan2018NIPS investigated the effectiveness of AT at the scale of ImageNet Deng2009CVPR , and proposed a logit pairing AT training method to tackle the tradeoff between robust accuracy and clean accuracy. Wong et al. Wong2020ICLR accelerated the training process using FGSM attack with random initialization instead of PGD attack Kurakin2017ICLR , and reached significantly lower cost. Xu et al. Xu2022NIPS proposed a novel attack method which can make a stronger perturbation to the input images, resulting in the robustness of models by AT using this attack method is improved. Li et al. Li2022CVPR revealed a link between fast growing gradient of examples and catastrophic overfitting during robust training, and proposed a subspace AT method to mitigate the overfitting and increase the robustness. Dabouei et al. Dabouei2022ECCV found the gradient norm in AT is higher than natural training, which hinders the training convergence of outer optimization of AT. And they proposed a gradient regularization method to improve the performance of AT.
@@
Robustness Enhancing without AT. Non-AT robustness enhancing techniques can be categorized into part-based models Li2023TPAMI ; Sitawarin2023ICLR , feature vector clustering Mustafa2021TPAMI ; Seo2023CVIU , adversarial margin maximization Yan2019TPAMI ; Guo2022TPAMI , etc. Li et al. Li2023TPAMI argued one reason that DCNNs are prone to be attacked is they are trained only on category labels, not on the part-based knowledge as humans do. They proposed an object recognition model, which first segments the parts of objects, scores the segmentations based on human knowledge, and final outputs the classification results based on the scores. This part-based model shows better robustness than classic recognition models across various attack settings. Sitawarin et al. Sitawarin2023ICLR also thought richer annotation information can help learn more robust features. They proposed a part segmentation model with a head classifier trained end-to-end. The model first segments objects into parts, and then makes predictions based on the parts. Mustafa et al. Mustafa2021TPAMI stated that making feature vectors in the same class closer and centroids of different classes more separable can enhance the robustness of DCNNs. They added PCL to the conventional loss function, and designed an auxiliary function to reduce the dimensions of convolutional feature maps. Seo et al. Seo2023CVIU proposed a training methodology that enhance the robustness of DCNNs through a constraint that applies a class-specific differentiation to the feature space. The training methodology results in feature representations with a small intra-class variance and large inter-class variances, and can improve the adversarial robustness notably. Yan et al. Yan2019TPAMI proposed an adversarial margin maximization method to improve DCNNs’ generalization ability. They employed the deepfool attack method Moosavi2016CVPR to compute the distance between an image sample and its decision boundary. This learning-based regularization can enhance the robustness of DCNNs as well.
@@
Note that, BEFB integrated models proposed in this paper can be easily combined with above-mentioned robustness enhancing techniques, such as AT Madry2018ICLR and PCL Mustafa2021TPAMI , and can achieve better classification accuracy than the original models.
@@ 3.1 Edge Detectors
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.2 Binary Edge Feature Branch
The threshold layer turns the output of the last Sobel layer to the binary edge features, which concatenated with the texture features learnt by the backbone are fed into fully connected layers for classification. Note that, the gradient of activation function of Eq. (5) is zero. To update the weights in BEFB, STE technique Bengio2013Arxiv ; Yang2022ICML ; Vanderschueren2023WACV is employed.
@@ 4 Experiments
In this section, we integrate BEFB into VGG16 Simonyan2014ICLR and ResNet34 He2016CVPR ; He2020TNNLS , respectively, and conduct the experiments on CIFAR-10 Krizhevsky2009 , MNIST Deng2012ISPM , SVHN Yuval2011NIPS , and TinyImageNet (TinyIN for short) Le2015Tiny datasets. We examine the effects of BEFB on model training, analyse obfuscated gradient effect Athalye2018ICML ; Yue2023USENIX ; Huang2022MSML by using different activation functions of the threshold layer to generate AEs, and compare the accuracy of BEFB integrated models when facing FGSM Goodfellow2015ICLR , PGD Kurakin2017ICLR , and C&W Carlini2017SP attacks with original ones. Furthermore, we combine the BEFB integrated models with two robustness enhancing techniques–AT Madry2018ICLR and PCL Mustafa2021TPAMI to evaluate the classification accuracy. All experiments are coded in Tensorflow with one TITAN XP GPU.
@@ 4.2 Effects of BEFB on model training
VGG16
99.17%
83.70%
18s
99.72%
99.40%
16s
99.68%
94.36%
18s
94.32%
33.84%
80s
VGG16-BEFB-multiple
99.53%
82.95%
18s
99.56%
99.46%
17s
99.35%
94.64%
18s
92.03%
30.01%
95s
ResNet34
99.32%
79.19%
48s
99.63%
99.40%
40s
99.76%
93.22%
50s
97.53%
30.80%
238s
ResNet34-BEFB-multiple
99.48%
74.58%
48s
99.38%
99.28%
42s
99.78%
93.41%
52s
97.98%
26.31%
250s
@@ 4.3 Analysis of obfuscated gradient effect
The activation function of the threshold layer is expressed by Eq. (5), whose gradient is zero. In order to update the weights in BEFB, STE technique Bengio2013Arxiv ; Yang2022ICML ; Vanderschueren2023WACV is adopted. But with respect to generating AEs, STE may bring obfuscated gradient effect Athalye2018ICML ; Yue2023USENIX ; Huang2022MSML , which means the yielded AEs are not powerful enough to deceive the models. Here, in addition to STE, zero gradient and gradient of a center-translated sigmoid activation function for threshold layer are also tested. Table 4 compares the classification accuracy of AEs generated by these three gradient computing strategies under FGSM attacks. Eq. (6) defines the center-translated sigmoid function. $x_{0}$ in general, is the threshold value for channels, illustrated in Eq. (5).
@@ 4.4 Performance comparison between BEFB integrated models and original models under attacks
VGG16
28.59%
1.93%
0.06
56.54%
3.63%
6.78
57.77%
16.84%
0.08
27.39%
4.03%
0.05
VGG16-BEFB-single
42.77%
8.01%
0.05
59.05%
5.83%
6.53
58.15%
17.98%
0.08
31.15%
8.74%
0.07
VGG16-BEFB-multiple
45.60%
8.05%
0.07
67.42%
15.53%
7.03 59.28%
19.48%
0.10
31.93%
9.18%
0.07
ResNet34
7.82%
0.47%
0.05
12.66%
0.76%
1.22
24.14%
2.49%
0.02
5.80%
0.00%
0.10
ResNet34-BEFB-single
10.31%
0.78%
0.06
13.33%
2.33%
1.50
23.95%
2.41%
0.02
5.99%
0.36%
0.11
ResNet34-BEFB-multiple
14.88%
1.01%
0.10
14.53%
2.95%
1.39
25.17%
2.98%
0.03
5.95%
0.95%
0.14

### v2

@@ 1 Introduction
It is well known that deep convolutional neural networks (DCNN for short) can be fooled by examples with small perturbations [1, 2, 3, 4, 5], which brings the potential hazards when applying DCNNs to the safety-critical applications, e.g., autonomous driving, airport security, and industry automation. Therefore, it is of great significance to improve the robustness of DCNNs. Fig. 1 shows several clean examples and their adversarial counterparts. Adversarial examples (AE for short) in Fig. 1(b) are with small noises which can induce DCNNs to make wrong decisions. However, these noises can be filtered by human eyes easily. Researches show that the principal way human beings recognize the objects is largely relying on the shape features [6, 7], and that’s why human beings are not susceptible to the noise imposed in Fig. 1(b). Inspired by this, it is natural to ask, is it possible to make the DCNNs more robust by learning the shape-like features?
@@
Fig. 2 shows the thresholded edge images of Fig. 1. From the figure, it can be observed that the thresholded edge images for Fig. 1(a) and Fig. 1(b) are almost the same. It prompts us that, enhancing the robustness of DCNNs can be achieved by using the binary edge features which can be seen as a kind of shape features. To this end, in this paper, four learnable edge detectors are designed and taken as layer kernels, which can be used to extract horizontal, vertical, positive diagonal, and negative diagonal edge features, respectively [8, 9, 10]. And based on the edge detectors, a binary edge feature branch (BEFB for short) is proposed, which is staked by multiple Sobel layers and one threshold layer. The Sobel layers employ the edge detectors as kernels, and the threshold layer turns the output of the last Sobel layer to the binary features. The binary features concatenated with the texture features learnt by any popular backbone are then fed into fully connected layers for classification. To deal with the zero gradient of threshold layer, which makes the weights in BEFB unable to update using the chain rule, STE technique [11, 12, 13] is employed. In addition, to take obfuscated gradient effect raised by [14, 15, 16] into consideration, zero gradient or one center-translated sigmoid activation function are also used for generating AEs. In the experiments, we integrate BEFB into VGG16 [17] and ResNet34 [18, 19], respectively, and conduct experiments on multiple datasets. The results demonstrate BEFB has no side effects on model training, and BEFB integrated models can achieve better accuracy under FGSM [1], PGD [20], and C&W [21] attacks compared to the original ones. Furthermore, we combine the BEFB integrated models with two popular robustness enhancing techniques–AT(abbreviation for adversarial training) [22] and PCL(abbreviation for prototype conformity loss) [23], respectively, and find BEFB integrated models can achieve better accuracy than the original models as well.
@@ 2 Related Work
Robustness Enhancing with AT. AT improves the robustness of the DCNNs by generating AEs as training samples during optimization. Tsipras et al. [24] found there exists a tradeoff between robustness and standard accuracy of the models generated by AT, due to the robust classifiers learning a different feature representation compared to the clean classifiers. Madry et al. [22] proposed an approximated solution framework to the optimization problems of AT, and found PGD-based AT can produce models defending themselves against the first-order adversaries. Kannan et al. [25] investigated the effectiveness of AT at the scale of ImageNet [26], and proposed a logit pairing AT training method to tackle the tradeoff between robust accuracy and clean accuracy. Wong et al. [27] accelerated the training process using FGSM attack with random initialization instead of PGD attack [20], and reached significantly lower cost. Xu et al. [28] proposed a novel attack method which can make a stronger perturbation to the input images, resulting in the robustness of models by AT using this attack method is improved. Li et al. [29] revealed a link between fast growing gradient of examples and catastrophic overfitting during robust training, and proposed a subspace AT method to mitigate the overfitting and increase the robustness. Dabouei et al. [30] found the gradient norm in AT is higher than natural training, which hinders the training convergence of outer optimization of AT. And they proposed a gradient regularization method to improve the performance of AT.
@@
Robustness Enhancing without AT. Non-AT robustness enhancing techniques can be categorized into part-based models [31, 32], feature vector clustering [23, 33], adversarial margin maximization [34, 35], etc. Li et al. [31] argued one reason that DCNNs are prone to be attacked is they are trained only on category labels, not on the part-based knowledge as humans do. They proposed an object recognition model, which first segments the parts of objects, scores the segmentations based on human knowledge, and final outputs the classification results based on the scores. This part-based model shows better robustness than classic recognition models across various attack settings. Sitawarin et al. [32] also thought richer annotation information can help learn more robust features. They proposed a part segmentation model with a head classifier trained end-to-end. The model first segments objects into parts, and then makes predictions based on the parts. Mustafa et al. [23] stated that making feature vectors in the same class closer and centroids of different classes more separable can enhance the robustness of DCNNs. They added PCL to the conventional loss function, and designed an auxiliary function to reduce the dimensions of convolutional feature maps. Seo et al. [33] proposed a training methodology that enhance the robustness of DCNNs through a constraint that applies a class-specific differentiation to the feature space. The training methodology results in feature representations with a small intra-class variance and large inter-class variances, and can improve the adversarial robustness notably. Yan et al. [34] proposed an adversarial margin maximization method to improve DCNNs’ generalization ability. They employed the deepfool attack method [36] to compute the distance between an image sample and its decision boundary. This learning-based regularization can enhance the robustness of DCNNs as well.
@@
Note that, BEFB integrated models proposed in this paper can be easily combined with above-mentioned robustness enhancing techniques, such as AT [22] and PCL [23], and can achieve better classification accuracy than the original models.
@@ 3.1 Edge Detectors
$w_{ij}\left\{\begin{array}[]{ll}\in\textrm{[0, 1]}&\mbox{if \emph{i} = 1, \emph{j} = 1, 2, 3},\\ =\textrm{0}&\mbox{if \emph{i} = 2, \emph{j} = 1, 2, 3},\\ \in\textrm{[-1, 0]}&\mbox{if \emph{i} = 3, \emph{j} = 1, 2, 3},\\ \end{array}\right.$ (1)
@@
$w_{ij}\left\{\begin{array}[]{ll}\in\textrm{[0, 1]}&\mbox{if \emph{j} = 1, \emph{i} = 1, 2, 3},\\ =\textrm{0}&\mbox{if \emph{j} = 2, \emph{i} = 1, 2, 3},\\ \in\textrm{[-1, 0]}&\mbox{if \emph{j} = 3, \emph{i} = 1, 2, 3},\\ \end{array}\right.$ (2)
@@
$w_{ij}\left\{\begin{array}[]{ll}\in\textrm{[0, 1]}&\mbox{if (\emph{i}, \emph{j}) $\in$ \{(1, 1), (1, 2), (2, 1) \}},\\ =\textrm{0}&\mbox{if (\emph{i}, \emph{j}) $\in$ \{(1, 3), (2, 2), (3, 1) \}},\\ \in\textrm{[-1, 0]}&\mbox{if (\emph{i}, \emph{j}) $\in$ \{(2, 3), (3, 2), (3, 3) \}},\\ \end{array}\right.$ (3)
@@
$w_{ij}\left\{\begin{array}[]{ll}\in\textrm{[0, 1]}&\mbox{if (\emph{i}, \emph{j}) $\in$ \{(1, 2), (1, 3), (2, 3) \}},\\ =\textrm{0}&\mbox{if (\emph{i}, \emph{j}) $\in$ \{(1, 1), (2, 2), (3, 3) \}},\\ \in\textrm{[-1, 0]}&\mbox{if (\emph{i}, \emph{j}) $\in$ \{(2, 1), (3, 1), (3, 2) \}},\\ \end{array}\right.$ (4)
@@ 3.2 Binary Edge Feature Branch
The threshold layer turns the output of the last Sobel layer to the binary edge features, which concatenated with the texture features learnt by the backbone are fed into fully connected layers for classification. Note that, the gradient of activation function of Eq. (5) is zero. To update the weights in BEFB, STE technique [11, 12, 13] is employed.
@@ 4 Experiments
In this section, we integrate BEFB into VGG16 [17] and ResNet34 [18, 19], respectively, and conduct the experiments on CIFAR-10 [37], MNIST [38], SVHN [39], and TinyImageNet (TinyIN for short) [40] datasets. We examine the effects of BEFB on model training, analyse obfuscated gradient effect [14, 15, 16] by using different activation functions of the threshold layer to generate AEs, and compare the accuracy of BEFB integrated models when facing FGSM [1], PGD [20], and C&W [21] attacks with original ones. Furthermore, we combine the BEFB integrated models with two robustness enhancing techniques–AT [22] and PCL [23] to evaluate the classification accuracy. All experiments are coded in Tensorflow with one TITAN XP GPU.
@@ 4.2 Effects of BEFB on model training
VGG16 99.17% 83.70% 18s 99.72% 99.40% 16s 99.68% 94.36% 18s 94.32% 33.84% 80s
VGG16-BEFB-multiple 99.53% 82.95% 18s 99.56% 99.46% 17s 99.35% 94.64% 18s 92.03% 30.01% 95s
ResNet34 99.32% 79.19% 48s 99.63% 99.40% 40s 99.76% 93.22% 50s 97.53% 30.80% 238s
ResNet34-BEFB-multiple 99.48% 74.58% 48s 99.38% 99.28% 42s 99.78% 93.41% 52s 97.98% 26.31% 250s
@@ 4.3 Analysis of obfuscated gradient effect
The activation function of the threshold layer is expressed by Eq. (5), whose gradient is zero. In order to update the weights in BEFB, STE technique [11, 12, 13] is adopted. But with respect to generating AEs, STE may bring obfuscated gradient effect [14, 15, 16], which means the yielded AEs are not powerful enough to deceive the models. Here, in addition to STE, zero gradient and gradient of a center-translated sigmoid activation function for threshold layer are also tested. Table 4 compares the classification accuracy of AEs generated by these three gradient computing strategies under FGSM attacks. Eq. (6) defines the center-translated sigmoid function. $x_{0}$ in general, is the threshold value for channels, illustrated in Eq. (5).
@@ 4.4 Performance comparison between BEFB integrated models and original models under attacks
VGG16 28.59% 1.93% 0.06 56.54% 3.63% 6.78 57.77% 16.84% 0.08 27.39% 4.03% 0.05
VGG16-BEFB-single 42.77% 8.01% 0.05 59.05% 5.83% 6.53 58.15% 17.98% 0.08 31.15% 8.74% 0.07
VGG16-BEFB-multiple 45.60% 8.05% 0.07 67.42% 15.53% 7.03 59.28% 19.48% 0.10 31.93% 9.18% 0.07
ResNet34 7.82% 0.47% 0.05 12.66% 0.76% 1.22 24.14% 2.49% 0.02 5.80% 0.00% 0.10
ResNet34-BEFB-single 10.31% 0.78% 0.06 13.33% 2.33% 1.50 23.95% 2.41% 0.02 5.99% 0.36% 0.11
ResNet34-BEFB-multiple 14.88% 1.01% 0.10 14.53% 2.95% 1.39 25.17% 2.98% 0.03 5.95% 0.95% 0.14